Andy is 5 years older than Chris in 3 years the sum of their ages will be 35. what are their ages
Answer by colliefan(242)

Andy is 5 years older than Chris in 3 years the sum of their ages will be 35. what are their ages
Algebraically:
Chris' age: x
Andy's age: x+5
In three years, both will be three years older so then:
Chris' age: x+3
Andy's age: x+5+3 or x+8
Then, you are given that in 3 years, the sum of the ages is 35 so:
(x+3) + (x+8) = 35
x+x+3+8 = 35
2x+11=35
2x+11-11=35-11
2x=24
2x/2=24/2
x = 12
Substitute this solution back in to make sure it is correct.
Chris' age = 12
Andy's age = 17